40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 4 20:42:14 2015, gautam, rana, Update, CDS, Checkout of the Wenzel dividers TEK00000.PNGTEK00001.PNGTEK00002.PNG
    Reply  Tue Sep 29 03:14:04 2015, gautam, Update, CDS, Frequency divider box IMG_0014.JPGIMG_0015.JPG
       Reply  Fri Oct 9 19:54:58 2015, gautam, Update, CDS, Frequency divider box - installation in 1X2 rack IMG_0027.JPGtime_seris_25MHz.pdf
          Reply  Mon Oct 12 17:04:02 2015, gautam, Update, CDS, Frequency divider box - further tests calibration.pdfsystematics.pdf
             Reply  Wed Oct 14 17:40:50 2015, gautam, Update, CDS, Frequency divider box - further tests time_series_input_signals.pdfcalibration_20151012.pdfsystematics_20151012.pdf
                Reply  Tue Oct 20 17:36:01 2015, gautam, Update, CDS, Frequency counting with moving average 
                   Reply  Fri Oct 23 18:36:48 2015, gautam, Update, CDS, Frequency counting - workable setup prepared Yscan.pdf
                      Reply  Fri Oct 23 19:27:19 2015, Koji, Update, CDS, Frequency counting - workable setup prepared 
                         Reply  Sat Oct 24 12:34:43 2015, gautam, Update, CDS, Frequency counting - workable setup prepared Yscan.pdfFrequency_readout.pdf
                      Reply  Thu Nov 5 03:04:13 2015, gautam, Update, CDS, Frequency counting - systematics and further changes Systematic_error.pdfsystematics_origin.pdf
Message ID: 11690     Entry time: Wed Oct 14 17:40:50 2015     In reply to: 11684     Reply to this: 11704
Author: gautam 
Type: Update 
Category: CDS 
Subject: Frequency divider box - further tests 

Summary:

I carried out some further diagnostics and found some ways in which I could optimize the zero-crossing-counting algorithm, such that the error in the measured frequency is now entirely within the expected range (due to a +-1 clock cycle error in the counting). We can now determine frequencies up to ~60 MHz with less than 1 MHz systematic error and <10 kHz statistical error (fluctuations after the 20 Hz lowpass). This should be sufficient for slow control of the end-laser temperatures.

Details: 

The conclusion from my earleir tests was that there was possibly an improvement that could be made to setting the thresholds for the Schmitt trigger stage in the model. In order to investigate this, I wanted to have a look at the 64K sampled raw input to the ADCs. Yesterday Eric helped me edit the appropriate .par file for viewing these channels for c1x03, and for an input frequency of 70MHz (after division, ~4.3 kHz square wave), the signal looked as expected (top left plot, attachment #1). This prompted me to check the counting algorithm again with the help of various test points I had setup within the model. I found that there was a tendency to under-count the number of clock-cycles between zero-crossings by more than 1 clock cycle, due to the way my code was organized. I fixed this and found that the performance improved dramatically, compared to my previous trials. With the revised counting algortihm, there was at most a +-1 clock cycle error in the counting, and the systematic error between the measured and requested RF frequencies is now completely accounted for taking this consideration into account. The origin of this residual error can be understood by looking at the top right plot in Attachment #1 - presumably because of the effects of some downsampling filter, the input signal to the Schmitt trigger isnt a clean square wave (even at 4kHz) - specifically, the time spent in the LOW and HIGH states of the Schmitt trigger can vary between successive zero crossings because of the shape of the input waveform. As a result, there can be a +-1 clock cycle error in the counting process. Attachment #2 shows this - the red and blue lines envelope the measured frequency for the whole range investigated: 10-70MHz. Attachment #3 shows the systematic error as a function of the requested frequency.

If there was some way to bypass the downsampling filter, perhaps the high-frequency performance could be improved a little. 

Attachment 1: time_series_input_signals.pdf  9 kB  | Hide | Hide all
time_series_input_signals.pdf
Attachment 2: calibration_20151012.pdf  23 kB  | Hide | Hide all
calibration_20151012.pdf
Attachment 3: systematics_20151012.pdf  18 kB  | Hide | Hide all
systematics_20151012.pdf
ELOG V3.1.3-